OpenStudy (joshoyen):

Solve each absolute value equation or inequality. 1. -3 + |m - 3| = 0

OpenStudy (solomonzelman):

1) Add 3 to both sides. 2) Use the fact that \(|\)f(x)\(|\)=r, can have two solutions, f(x)=\(\pm\)r

OpenStudy (joshoyen):

m-3=3 ?

OpenStudy (solomonzelman):

yes, or m-3=-3

OpenStudy (solomonzelman):

if, n-3=3, then n=?

OpenStudy (solomonzelman):

(I mean, m - not n, but you get the point)

OpenStudy (joshoyen):

m=6

OpenStudy (solomonzelman):

yes, good

OpenStudy (solomonzelman):

and if m-3=-3 then, m=?

OpenStudy (joshoyen):

m=0

OpenStudy (solomonzelman):

So, your two answers/solutions are: m=0 m=6

OpenStudy (joshoyen):

ohhh okay thanks!

OpenStudy (solomonzelman):

Anytime!

OpenStudy (joshoyen):

Wait how would I do 1 + |5x - 6| = 47

OpenStudy (solomonzelman):

1) Subtract 1 from both side. This will get you: \(|\)5x-6\(|\)=3 Then, your two solutions are; 5x-6=3 and 5x-6=-3

OpenStudy (solomonzelman):

solve each equation individually and that will give you the two answers.

OpenStudy (joshoyen):

how did you get 3

OpenStudy (solomonzelman):

oh, 47, I read 4... apologize

OpenStudy (joshoyen):

oh haha

OpenStudy (solomonzelman):

CORRECTION *** 1) Subtract 1 from both sides. This will get you: \(|\)5x-6\(|\)=46 Then, your two solutions are; 5x-6=46 and 5x-6=-46

OpenStudy (solomonzelman):

then, solve each equation for x....

OpenStudy (joshoyen):

alright i see now, thank you!

OpenStudy (solomonzelman):

You are welcome!

OpenStudy (joshoyen):

@SolomonZelman Hey could you help me on this otherp roblem it's |-4+x| + 7 < 18

OpenStudy (solomonzelman):

You want to isolate absolute value by itself, so what would you do to accomplish this?

OpenStudy (joshoyen):

hm i'm not sure, i tried subtracting 7 from both sides but that didnt work

OpenStudy (solomonzelman):

subtracting 7 is right .... what did you get after subtracting 7?

OpenStudy (joshoyen):

oh wait nvm I didn't even subtract 7, lmao i added 7.. wht was i thinking

OpenStudy (joshoyen):

okay i got it now

OpenStudy (joshoyen):

thanks lol

OpenStudy (solomonzelman):

sure\( ; ) \)